OpenStudy (anonymous):

area of a rectangular object which has a length of inches and a width of 4sqrt2 and a width of 6sqrt6

OpenStudy (bahrom7893):

A = 4sqrt2 * 6sqrt6 = 24 sqrt12 = 24 * sqrt(4*3) = 24*2 sqrt3 = 48sqrt3